1. A battery with a terminal voltage of 12V is connected to a circuit consisting of four 10 and one 5 resistor all in series. Assume the battery has negligible internal resistance. Calculate the current through each resistor.




2. Calculate the RC time constant, τ of a circuit having a resistance of 15 MΩ and 3 010 𝜇𝐹.




3. An RC circuit have a 9 V source and lies at a 54-seconds of time constant. Then, find the voltage across the Capacitor after 153.6 seconds of charging.

1
Expert's answer
2022-04-28T16:41:34-0400
"I=\frac{V}{R}=\frac{12}{10+5}=0.8\:\rm A"

2.

"\tau=RC=15*10^6*3010*10^{-6}=45150\;\rm s"